# ZMP-HH-25/5

### v1

@@ C.2 The Clemens–Schmid exact Sequence
$0\to H^{3}(\mathcal{V})\stackrel{{\scriptstyle\iota^{*}}}{{\to}}H^{3}(V_{z})\stackrel{{\scriptstyle N}}{{\to}}H^{3}(V_{z})\to H_{3}(\mathcal{V})\to\cdots\,,$ (C.11)
@@
$0\to H^{3}(V_{0})\stackrel{{\scriptstyle\iota^{*}}}{{\to}}H^{3}(V_{z})\cap\mathrm{ker}\,N\stackrel{{\scriptstyle N}}{{\to}}0\,,$ (C.12)
@@
$\cong\begin{cases}H^{3}(X_{1})\oplus H^{3}(X_{2})\,,&\ell=3\,,\\ H^{2}(Z)/(\mathrm{im}(i^{*})+\mathrm{im}(j^{*}))\,,&\ell=2\,.\end{cases}$
This gives us a full geometric understanding of the lower half of the corresponding Hodge–Deligne diamond depicted in Figure -958, which is sufficient for our purposes. [^47] The attentive reader may wonder whether there is also a direct geometric interpretation of the space $\mathrm{Gr}_{4}$. The answer is yes. In terms of the weight filtration, one can identify $W_{4}=H^{3}_{\mathrm{log}}(V_{0})$, where the latter denotes the logarithmic de Rham cohomology group associated to $V_{0}$. The reason for including such logarithmic forms is that the fourfold $\mathcal{V}$ is non-compact. This is explained in detail in Section 5 of [106].

### v2

@@ C.2 The Clemens–Schmid exact Sequence
$0\to H^{3}(\mathcal{V})\mathrel{{\mathop{\to}\limits^{\iota^{*}}}}H^{3}(V_{z})\mathrel{{\mathop{\to}\limits^{N}}}H^{3}(V_{z})\to H_{3}(\mathcal{V})\to\cdots\,$, (C.11)
@@
$0\to H^{3}(V_{0})\mathrel{{\mathop{\to}\limits^{\iota^{*}}}}H^{3}(V_{z})\cap\mathrm{ker}\,N\mathrel{{\mathop{\to}\limits^{N}}}0\,$, (C.12)
@@
[…]
This gives us a full geometric understanding of the lower half of the corresponding Hodge–Deligne diamond depicted in Figure 2, which is sufficient for our purposes. [^48] The attentive reader may wonder whether there is also a direct geometric interpretation of the space $\mathrm{Gr}_{4}$. The answer is yes. In terms of the weight filtration, one can identify $W_{4}=H^{3}_{\mathrm{log}}(V_{0})$, where the latter denotes the logarithmic de Rham cohomology group associated to $V_{0}$. The reason for including such logarithmic forms is that the fourfold $\mathcal{V}$ is non-compact. This is explained in detail in Section 5 of [108].